HUBBLE SPACE TELESCOPE DAILY REPORT #2519 PERIOD COVERED: 0000Z (UTC) 12/13/99 - 0000Z (UTC) 12/14/99 Daily Status Report as of 348/0000Z 1.0 OBSERVATIONS SCHEDULED AND ACCOMPLISHED: Since HST remains in Zero Gyro Sun Point (ZGSP), no science observations were performed. 2.0 FLIGHT OPERATIONS SUMMARY: 2.1 Guide Star Acquisitions: No activity 2.2 FHST Updates: No activity 2.3 Operations Notes: The HST continues in ZGSP. Per an operations request, CCC 5 K2 was set to level 4 at 347/1518Z. The time-based recharge ratio charge/discharge buffers were reset at 347/1522Z as directed by Operations Note #661. Operations Note #686 was utilized to adjust the CCC K5 V/T level at 347/1532Z. A recycle of the CCS "A", "B", and "G" strings was requested in preparation for Servicing Mission 3A was requested. It was noticed that the FEP and the backbone were receiving data but the NTs were not. HSTAR 7402 was written. Another recycle of all three strings was successfully accomplished at 347/2000Z. A TTR was written for T-1 line dropouts at 348/0841Z. The problem was corrected by a line switch. The resultant missing data will be recovered by a requested line outage recorder playback. 3.0 SIGNIFICANT FORTHCOMING EVENTS: Servicing Mission 3A, currently scheduled for December 16, 1999. /CAW